OpenStudy (anonymous):

Explain how you would find the exact value of 3 / sqrt 8.

OpenStudy (jack1):

sqrt 8 = 8^-1/2

OpenStudy (jack1):

=(sqrt 2) / (4)

OpenStudy (jack1):

so \[3/\sqrt{8} = (3\times \sqrt{2})/4\]